Backlinks to BCU in all Webs (Search ARGOSPublic Web only)

Results from ARGOSPublic web retrieved at 19:38 (GMT)

BcuNumber1
Known history When Where Who What 201204xx OAA BCU used by HD SW group to develop the software is shipped to Arcetri 20120412 OAA OAA ...
BcuNumber2
Known history When Where Who What 20130215 delivered to Arcetri in the DX rack delivery 20130220 OAA LB Flashed new Logic version to s...
BcuNumber3
Known history When Where Who What 20130716 OAA BcuNumber3 received from MPE 20130730 OAA MB Exchanged with BcuNumber1 in LgswRack2 ...
CablePiezoBCU0
Known history When Where Who What 20130605 OAA MB Installed in LgswUnitSx in the BlueChannel. 20150618 OAA MB Removed from LgswUnitSx ...
CablePiezoBCU1
Known history When Where Who What 2014xxxx LBT RB Shipped to LBT as spare cable Know issues When What 20130305 description of issu...
CablePiezoBCU10
Known history When Where Who What 20150604 OAA MB It is on the shelf in the ExIbis lab 20150618 OAA MB Shipped to LBT as spare Know...
CablePiezoBCU2
Known history When Where Who What 2014xxxx LBT RB Shipped to LBT as spare cable Know issues When What description of issue Main...
CablePiezoBCU3
Known history When Where Who What 20150604 OAA MB It is on the shelf in the ExIbis lab 20150818 LBT MB Installed on LgswUnitSx in the ...
CablePiezoBCU4
Known history When Where Who What 20130215 OAA Installed in LgswUnitDx in the RedChannel. Used to connect PiezoS334616 to BcuNumber2 Know iss...
CablePiezoBCU5
Known history When Where Who What 20130215 OAA Installed in LgswUnitDx in the YellowChannel. Used to connect PiezoS334617 to BcuNumber2 Know ...
CablePiezoBCU6
Known history When Where Who What 20130215 OAA Installed in LgswUnitDx in the BlueChannel. Used to connect PiezoS334615 to BcuNumber2 Know is...
CablePiezoBCU7
Known history When Where Who What 20130605 OAA MB Installed in LgswUnitSx in the YellowChannel. Know issues When What description ...
CablePiezoBCU8
Known history When Where Who What 20150605 OAA MB Installed in LgswUnitSx in the RedChannel. Know issues When What description of ...
CablePiezoBcu
Description Cables used to connect the PI S334 to the BCU. Each piezo head has 5 cables terminated with LEMO connectors: 3 for the HV commands (Axis 1, Axis 2, 10...
DayTime20140513
We install the external laser source on the Red arm. 10:00 Micheal Wagner disconnects the accelerometer board on the DX ASM. Acceptance test positions for the Red...
DayTime20141102
DX LGSW setup 9:00 a.m.: we start up the LGSW. There is a problem with the state machine: in the powered off state the Thors are off and when you restart, passing...
DayTime20141103
9:00: dichroic CalUnit in M3 758 580 0 25.9 Hexapod coma free movements: X= 0.982mm, Ry= 102" Y= 0.982mm, Rx= 102" Slided Red entrance window by...
DayTime20141104
Calibrations 9:30: start of operations HEXA 2.03 7.25 4.47 31.74 187 0 We have a problem with BCU: when we change subapertures the HVC goes to ...
DayTime20141106
9:24: system fully started, LGSW, ADSEC, FLAO, IRTC are ready. On the FLAO we load the board setup 2014 05 16 ARGOS cal onaxis. Board stage mm X 70 ...
DayTime20141126
8:00: we start setting up the system for calibrations. HEX X Y Z Rx Ry 3.02 6.27 4.35 324 155 Acquired new DARK: 20141126_160957 We...
DayTime20141127
8:30: start of calibration unit activity, spots are already in LGSW HEXA X Y Z Rx Ry 5.02 7.27 4.38 365 135 We close the loop witho...
DayTime20141128
8:30: we start setting up the system for calibrations 9:30: ready to close the loop on ARGOS we acquire a Board setup: 28 Nov 14_argos_calibunit with FLAO BIN4 PB...
DayTime20141130
We connect the BCU power supply line to the Agilent oscilloscope. 18:00: we start operation swith calibration unit monitoring the BCU voltage. 20:00: after repea...
DayTime20150430
Best focus on LUCI2 15:00: switch BCU swap: SX one is now in the DX ASM 17:40: cal unit in, we close on LGSW only HEXA 2.94 8.18 2.08 150 12 Boa...
DayTime20151210
PBL: we have found a problem with slopes reordering... We close a loop with LGSW only REC: 20151209_213200 w/o disturb. TN AO Gain Res. WF nm rms AO rec...
DayTime20151212
Test of DX LGSW with CalUnit and rigid secondary Hexapod settings with rigid secondary: Right PSF 2ndary X Y Z Rx Ry Rz 4.75 8.30 5.15 4.9...
DayTime20160512
5:30: Gustavo deploys the arm, we have pbl starting AdSecSx: it periodically faild while loading the DSP program. We try many times to power cicle. LB finally can...
DayTime20160515
Fast guiding check w/ oscilloscope 11:30: Tommaso and Sebastian go on the telescope to check the Fast Guiding delay. We connect an oscilloscope after the 10x divi...
DayTime20161012
Visual Inspection 11:00: telescope back to Zenith, we climb and check the racks: fine. The pressure gauge on SX PNCCD had low battery. We pumped it anyway up to 0...
DayTime20161013
LgswUnitSx AO calibration 6:45: start of operations, we try to power up from arbitrator. SX BCU fails to boot twice, then we power cicle from Rack I/O and it work...
DayTime20161016
LgswUnitDx work 8:30: WG and MR go to check the cal unit illumination patter. It seems ok, maybe an alignment problem? 10:20: BcuNumber3 fails to connect. We must...
DayTime20170301
14:00: We switch on DX side. DX LGSW pulse generator is working, WFS camera and BCU are working at 970 Hz. 15:00: We pump the SX WFS camera, pressure was 90mbar o...
DiaryRun12
Log * List of action items for the run: 201605_ToDoList_SR_.xlsx When What Log Notes 20160503 Arrival at telescope of mech team 20160510 ...
DiaryRun2
Who When M. Bonaglia 20140506 20140520 L. Busoni 20140506 20140520 T. Mazzoni 20140506 20140520 M. Kulas 20140508 20140520 J. ...
DiaryRun3
Who From To T. Mazzoni 20141025 20141103 J. Ziegleder 20141028 201411?? G. Orban de Xivry 20141028 201411?? W. Raab 20141028 201...
DiaryRun4
Plan * Task_ for_November_ run_2014.docx: Task_ for_November_ run_2014.docx * ARGOS_Nov28_To_Dec2_2014.txt: TargetListRun4 Pictures Activity summary When ...
DiaryRun6
Plan Hi all, Please find below a summary of the activities concerning the installation of the ARGOS SX dichoric foreseen for the late April '15 run. 22/04 (Wed): ...
DiaryTecRun
Main.SebastianRabien 12 Aug 2016 Julian and SR on the mountain. Actions performed: * The new remote power is installed and tested. New engineering gui installe...
EngPanel
LBT EngPanel LBT EngPanel is a tool developed by Microgate to configure and connect to a BCU unit. The attached software runs on Windows machines, and require the...
HowToAdjustBcuPowerSupply
* OBSOLETE: THE ARGOS BCU POWER SUPPLIES HAVE BEEN UPGRADED BY MICROGATE IN 2016 * Informations about how to read the BCU power supply voltage directly from the ...
HowToInstallARGOSFiberSplitter
* * IMPORTANT: This configuration has been removed. Final solution uses a fiber switch. It is left here for reference only. (GR/2017 02 10) See HowToUseTheFiberSw...
HowToPowerUP
Pre requisits: * Organize telescope access with Telescope Manager and mail to telescopework * ARGOS chiller must be running (HowToStartChiller) * A...
HowToRecoverArgos
How to recover ARGOS Procedure to recover from BCU crash 1 Set LAS and LAN in simulation (Arbitrator GUI, Tab: Expert Commissioning) 1 Open LGSW engineeri...
HowToSetupARGOSForCalibrations
How To Setup ARGOS for AO calibrations Pre requisits: * Organize telescope access with Telescope Manager and mail to telescopework * Telescope must be at...
HowToStartSoftware
How to start Software There are multiple ways to start up the software and depending on situation one or the other is more appropriate. The ARGOS software has als...
LabGUI
LabGUI This is the main engineering GUI that allows to control the full LgswSystem hardware. #StartLabGUI How to start the LabGUI i. Login on the workstation:...
MicrogateSrl
Microgate srl is the company that produced the BCU. Contacts Web: www.microgate.com Mario Andrighettoni Roberto Biasi
NightTime20141129
The list of targets and LUCI images is here: raw_20141129.txt 21:00: we start the telescope collimation and we resume the tubone. PatCam X Y Blue 172 ...
NightTime20151214
Very snowy day. Telescope is closed. We pull the LgswUnitSx 10mm forward (towards dichroic). LGSW focus now corresponds to Z=45mm. Debug of BCU communication, we ...
NightTime20160312
19:30: telescope open, flats taken, we notice a strange pattern on LUCI1 detector, probably it is due to consensation on LUCI entrance window. NGC2419 20:20: pre...
NightTime20160313
16:00: calibration of Hexapod to Patrol Cameras on the SX system: REC= 20160313_220000. This will be used for the fast laser guide w/ HVC green channel... Truth s...
NightTime20160515
18:00: back to SX, we setup for the night We start testing the new telescope offset procedure. We have a first try and the AOS hangs: Mar May 16 03:05:56 DX syste...
NightTime20160516
20:00: we are on sky with lasers 20:30: FLAO fails, CCD39 was off at 3:29:49, we had an error in the FYSOM state machine for OPT.LOOPBUF NOMAD 1499_0210548 * ...
NightTime20160517
DT: training of XY on SW and HW LgswUnitSx calibrations 20:30: we deploy the cal unit and we try to do SX calibrations SX HEXA 2.79 0.32 0.78 208 3...
NightTime20160518
BCU SX debugging 15:00: we start to check the behaviour of the SX BCU. We notice that the main issue is when the Pyramid slopes are enabled. On DX we measure slop...
NightTime20160519
20:00: problems w/ Sx AdSec, it goes in panic mode at each restart. SDSS J13432.85 415503 20:39: AdSec started. Propagate lasers. TN AO Gain AO CL TT Res...
NightTime20161020
19:00 22:00: time allocated to LUCI/AO G111.25 * Problem with the sign in the stages directions signs, solved * Pointing reference position for FLAO is 7...
NightTime20161022
10:00 LAT/05:00 UT Handover from LUCI/AO * Lorenzo Gustavo where switching fibers. * SX dichroic got stuck while authorization. * Authorization failed ...
NightTime20170503
19:30: getting ready to open... LEGA C 1 20:00: preset sent, longslit observation. NGS magnitude about 14. * PBL when 0;0 offset is sent on DX everything go...
NightTime20170504
19:45: collimating... Lasers are on sky. LEGA C 1 20:00: longslit observation. NGS magnitude 14. SX system closes loop w/o any problem on DX we get many instead:...
NightTime20170505
PSF wobble on LUCI1 We close the loop on FLAO, we put in N30 mask and we spin the rotator w/ and w/o the TTOffload LUT: * SX_ROTANG_TT_ON.txt: SX_ROTANG_TT_ON....
ThingsThatIDontUnderstand0001
Device The BCU #1 that was first delivered to Florence (and was before in Heidelberg for SW development). It is without ADC board, and was always affected by prob...
ThingsThatIDontUnderstand0002
Device We had this issue on BCU #1 and also on BCU #2 Description Quando si configura le 2 schede HVC, qualcosa va storto e la scheda di sotto non si configura...
ThingsThatIDontUnderstand0005
Device This issues affects BCU #2 and maybe other BCUs since it seems to be a firmware issue Occurrence When Notes 20130312 First observed. Repeatable,...
ThingsThatIDontUnderstand0008
Occurence When notes 20130820 Description We got an error with opcode 202 on BcuNumber3 when writing into DSP memory during frame grabbing. This could...
ThingsThatIDontUnderstand0009
Occurence When notes 20121113 Clearly reproducible on BcuNumber3 on LgswSystemSx Description After a "fresh" reset of the BCU the initialization goes...
VirtualBox
Both LGSW workstations WorkStation1, WorkStation2 have a Windows XP OS running on Oracle VM virtualbox. On Windows XP the following software is installed: * Mi...
WebStatistics
%IF{ "istopic '.DefaultWebStatistics'" then="$percentINCLUDE{$quot.DefaultWebStatistics$quot}$percent" else="$percentINCLUDE{$quot.DefaultWebStatistics$quot}$perc...
Number of topics: 68

No backlinks in public webs
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding AOWiki? Send feedback